OpenStudy (strangeviolets):

Which of the following is the equation of a line parallel to the line y=4x+1 and passes through the point (5,1) a. 4x+y=19 b. 4x-y=19 c. 4x+y=-19 d. -4x-y=19

OpenStudy (photon336):

parallel equations have the same slope

OpenStudy (photon336):

just use the point slope formula \[m(x-x_{o}) = (y-y_{o})\] \[4(x-5) = (y-1)\] do you know how you would simplify this?